OpenStudy (anonymous):

Transform this expression to radical form x^1/2y^23.Please show the complete solution. Transform this expression to radical form x^1/2y^23.Please show the complete solution. @Mathematics

OpenStudy (mathstudent55):

Do you mean \[x ^{\frac{ 1 }{ 2 }}y ^{\frac{ 2 }{ 3 }}\] ?

OpenStudy (mathstudent55):

\[a ^{\frac{ m }{ n }} = \sqrt[n]{x ^{m}}\]